If the probability of the Seattle Mariners winning the World Series is 2/30 What are the odds of the Mariners winning the World Series? : What is the probability of the Mariners not winning? What are the odds of the Mariners not winning?
Answer #1

Answer

odds of the Mariners winning the World Series = (number of ways Seattle Mariners winning the World Series)/(number of ways Seattle Mariners not winning the World Series)

= 2/(30-2)

= 2/28

= 1/14

So, odds of winning = 1: 14

Probability of the Mariners not winning = 1 - Probability of the Mariners winning

= 1 - 2/30

= (30-2)/30

= 28/30

odds of the Mariners not winning the World Series = (number of ways Seattle Mariners not winning the World Series)/(number of ways Seattle Mariners winning the World Series)

= 28/2

= 14/1

= 14:1